OpenStudy (anonymous):

Find the points y=-2x+7 and y=1/3x (_____,_____)

OpenStudy (anonymous):

please can yall help me

OpenStudy (anonymous):

substitution brah

OpenStudy (anonymous):

y=-2x+7 [1] y=1/3x [2] plug in [2] into [1] so 1/3x=-2x+7 1/3x+2x=7 1/3x+6/3x=7 7/3x=7 x=7/(7/3) x=?

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

please hurry its going to log me out